In 2.97 minutes, a ski lift raises four skiers at constant speed to a height of 131 m. The average mass of each skier is 69.6 kg. What is the average power provided by the tension in the cable pulling the lift?
Answer #1

Pavg = U/t

U = mgh

t = 2.97 min = 2.97 *60 s

Pavg = 69.6 * 9.8 * 131 / (2.97 * 60)

Pavg = 501.42 W